I don't understand how to do this question:
Write the equation of the line that passes through the given point and is parallel to the given line. Give the answer in slope-intercept form. (-7,4) y+5x =-11
y=
thank you!

Found 2 solutions by nerdybill, dabanfield:
Answer by nerdybill(7384)

Write the equation of the line that passes through the given point and is parallel to the given line. Give the answer in slope-intercept form. (-7,4) y+5x =-11
.
First, determine the slope of the "given line":
y+5x =-11
y = -5x-11
Because the above is in the "slope-intercept" form we see that the slope is -5.
.
For two lines to be parallel, they must have the same slope.
.
Therefore, our new line has a slope of -5 and passes through (-7,4).
.
Use the "point-slope" form:
y-y1 = m(x-x1)
y-4 = -5(x-(-7))
y-4 = -5(x+7)
y-4 = -5x-35
y = -5x-31 (this is what they're looking for)

Answer by dabanfield(803)

Write the equation of the line that passes through the given point and is parallel to the given line. Give the answer in slope-intercept form. (-7,4) y+5x =-11
The slope-intercept form of a line is:
y = m*x + b where m is the slope of the line and b is the value of y where the line crosses the y axis (that is when x=0).
If we take the line y + 5*x = -11 and rewrite it in slope-intercept form we have:
y = -5x - 11
In this form we can see the slope of the line is -5.
Any line parallel to this line must also have a slope of -5 so the equation of the parallel line looks like:
y = -5x + b
We still need to find the value of b. Since we are told the point (-7,4) is on the line we can substitute -7 for x and 4 for y in the equation above and get:
4 = -5*-7 + b
b = 39
So the equation of the parallel line is:
y = -5x + 39